OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. (2 points) u = <2, -4>, v = <3, -8> 3.0° 6.0° -7.0° 16.0°

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

find the dot product first

OpenStudy (anonymous):

6 and 32

OpenStudy (misty1212):

lol the dot product is a number, not a choice of numbers

OpenStudy (anonymous):

38

OpenStudy (misty1212):

ok then find the absolute value of each vector

OpenStudy (anonymous):

how

OpenStudy (misty1212):

\[||<a,b>||=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

but u just said it was one # XD

OpenStudy (misty1212):

oh my

OpenStudy (misty1212):

we have a long way to go don't we?

OpenStudy (anonymous):

i mean it depends how much time u want to waste with me ill forget it tomorrow lol

OpenStudy (misty1212):

\[\cos(\theta)=\frac{u\cdot v}{|u|\times |v|}\]

OpenStudy (anonymous):

what is that

OpenStudy (anonymous):

cosine theta and then i draw a blank lol

OpenStudy (misty1212):

the numerator is the dot product, you got 38

OpenStudy (misty1212):

the denominator is the product of the absolute value of the vectors

OpenStudy (anonymous):

ok can u give me the numerator please

OpenStudy (misty1212):

\[|<2,-4>=\sqrt{2^2+4^2}=\sqrt{20}\]

OpenStudy (misty1212):

do the same for \(v\)

OpenStudy (anonymous):

i still dont understand lol

OpenStudy (anonymous):

like i get what u did to get here but what is v

OpenStudy (misty1212):

the other vector

OpenStudy (anonymous):

what is the other vector

OpenStudy (misty1212):

you wrote it , not me

OpenStudy (anonymous):

i have 38 and i dont knwo what is after

OpenStudy (anonymous):

please i want to go to sleep can i please just have the asnwer

OpenStudy (misty1212):

38 is the numerator you need the denominator too

OpenStudy (anonymous):

im gonna cry please i have like 1 minute left on my quiz

OpenStudy (misty1212):

this is what you are going to use \[\cos(\theta)=\frac{u\cdot v}{|u|\times |v|}\] we got \(u\cdot v=38\)

OpenStudy (anonymous):

i need a good grade if not ill get beat

OpenStudy (anonymous):

jk that wont happen lol

OpenStudy (anonymous):

dont u get the same thing as the denominator???

OpenStudy (misty1212):

looks like \(6^\circ\)

OpenStudy (anonymous):

so thats the asnwer?!?!?!